OpenStudy (anonymous):

2 |x- 2| greater than or equal to 3

jimthompson5910 (jim_thompson5910):

2 |x- 2| >= 3 |x- 2| >= 3/2 x- 2 >= 3/2 or x-2 <= -3/2 x >= 3/2+2 or x <= -3/2+2 x >= 7/2 or x <= 1/2